Show that the set of fuctions y1=e^-2x , y2=e^6x is linearly independent on (minus infinity, posi...

show that the set of fuctions y1=e^-2x , y2=e^6x is linearly independent on (minus infinity, positive infinity)
Answer #1

Solution:

Consider \small ay_{1}+by_{2}=0

\small ae^{-2x}+be^{6x}=0

Put \small x=0

\small \therefore a+b=0.........\left ( i \right )

Differentiate \small ae^{-2x}+be^{6x}=0    with respect to \small x .

\small \therefore -2ae^{-2x}+6be^{6x}=0

Put \small x=0

\small \therefore -2a+6b=0\Rightarrow -a+3b=0...............\left ( ii \right )

Solving \small \left ( i \right )    and \small \left ( ii \right )    simultaneously we get,

\small a=b=0

Thus, \small y_{1}    and \small y_{2} are linearly independent   on \small \left ( -\infty ,\infty \right ) .